I have to solve this algebraically, not by using a graph: A parabola has a vertex (8, 3) and one x intercept of 5. Determine the coordinates of the other x- intercept.
Answer by ewatrrr(10682)

 
Hi,
solving algebraically: Parabola with V(8,3) and passing thru (5,0)
y = a(x -8)^2 + 3
0 = a(5-8)^2 + 3
-3/9 = a
y = (-1/3)(x-8)^2 + 3
0 = (-1/3)(x-8)^2 + 3
-3 = (-1/3)(x-8)^2
(x-8)^2 = 9
(x-8) = ± 3
x = 5 and x = 11
(11,0) are the coordinates of the other x- intercept.